OpenStudy (jjenk):

What are the zeros of the polynomial function: f(x) = (x - 3)(x + 1)(x - 5)?

OpenStudy (whpalmer4):

When you write the polynomial in factored form like that, you can just read out the zeros, because they are the solutions to factor1=0 factor2=0 factor3=0 etc. so here \[x-3=0\]\[x+1=0\]\[x-5=0\]

OpenStudy (whpalmer4):

If you substitute any of the values you find by solving those trivial equations into the original polynomial, you'll see that the product is 0.